Find the volume of the solid formed by revolving the region bounded by y=(x-2)² and y=x about the y -axis.



1
Expert's answer
2021-07-21T12:55:16-0400

Solution;

Using the Shell Method,

Find the intersection points of the two functions,y= x and y=(x-2)2

At point of intersection,the functions are equal,

x=(x-2)2

x=x2-4x+4

x2-5x+4=0

Which is a quadratic equation,

Solving

"\frac{5_-^+\sqrt{5^2-(4×1×4)}}{2}"

x=4 and x=1

Volume by shell method;

V="\int_a^b2πx(g(x)-f(x))"dx

V="\int_1^42πx(x-(x-2)^2)dx"

V="\int_1^42πx(x-(x^2-4x+4))dx"

V="\int_1^42π(5x^2-x^3-4x)dx"

V="[2π(\frac{5x^3}{3}-\frac{x^4}{4}-2x^2)]_1^4"

V=2π×11.243

V=70.64185

Answer:

70.64 cubic units.